# Positive operator-valued noncommutative polynomials are squares

### v1

@@ 5.1. The positive block matrix $S$
For $v,w\in\left\langle x\right\rangle_{d+1}$ define the finite rank operators
@@
Denote by $S$ the $N(d+1)\times N(d+1)$ block operator matrix whose $(v,w)$ entry is $S_{w^{*}v}.$ If $p=\sum_{w\in\langle x\rangle_{d+1}}P_{w}\,w$ and $q=\sum_{v\in\langle x\rangle_{d+1}}Q_{v}\,v$, a direct computation gives the block–trace identity
@@
Let $\mathcal{H}^{N(d+1)}:=\bigoplus\limits_{w\in\langle x\rangle_{d+1}}\mathcal{H}.$ Define the row operator $P:\mathcal{H}^{N(d+1)}\to\mathcal{H}$ by
@@
To prove that $S$ is a positive operator, it remains to show that $S$ is self-adjoint. For $T\succeq 0,$ equation (5.2) says $\operatorname{Tr}(ST)$ is real. Hence,
$\operatorname{Tr}(S^{*}T)\ =\ \operatorname{Tr}(TS^{*})\ =\ \operatorname{Tr}((ST)^{*})\ =\ \overline{\operatorname{Tr}(ST)}\ =\ \operatorname{Tr}(ST).$
$\operatorname{Tr}(S^{*}T)\ =\ \operatorname{Tr}(ST)\quad\text{for all $T\in\mathcal{B}(\mathcal{H}^{N(d+1)})$ }.$
Hence $S^{*}=S$ and $S\succeq 0.$ Because $J$ and $\langle x\rangle_{d+1}$ are finite, $S$ is finite rank.
@@ 5.2. An Auxiliary Hilbert space $\mathcal{M}$
We construct a finite-dimensional Hilbert space $\mathcal{M}$ from the psd sesquilinear form induced by the psd operator $S.$ The Hilbert space $\mathcal{E}$ is constructed as a subspace of $\mathcal{M}$ in Step 4, Subsection 5.4.
@@
$\mathcal{N}\ :=\ \{z\in V:\langle z,z\rangle_{V}=0\}$
@@
$\mathcal{M}\ :=\ V/\mathcal{N}.$
Since $S$ has finite rank, $\mathcal{M}$ is finite dimensional (hence complete).
@@ 5.3. The coordinate maps $\Phi(w)$
For each $w\in\langle x\rangle_{d+1}$ define
$\Phi(w):\mathcal{H}\to\mathcal{M},\qquad\Phi(w)\xi\ =\ \big[(\delta_{u,w}\xi)_{u\in\langle x\rangle_{d+1}}\,\big],$
@@ 5.4. The Hilbert space $\mathcal{E}$ and the operator tuple $Y$
$L_{x_{i}}\big(\Phi(w)\xi\big)\ =\ \Phi(x_{i}w)\,\xi\qquad(w\in\left\langle x\right\rangle_{d},\ \xi\in\mathcal{H}).$ (5.4)
@@ 5.4.1. The $L_{x_{j}}$ are well-defined
Since the vectors $\Phi(v)\eta$ span $\mathcal{M}$, we have $L_{x_{i}}z\in\mathcal{N},$ proving that $L_{x_{i}}$ is well-defined. Let $P_{\mathcal{E}}$ be the orthogonal projection of $\mathcal{M}$ onto $\mathcal{E}.$ For $i=1,\dots,g,$ define $Y_{i}:\mathcal{E}\to\mathcal{E}$ by
@@ 5.5. The representing vector and evaluation
We construct the representing vector $\gamma$ and complete the proof.
Let $\mathrm{vec}:HS(\mathcal{H},\mathcal{E})=B(\mathcal{H},\mathcal{E})\to\mathcal{H}\otimes\mathcal{E}$ denote the vectorization map defined on the space of Hilbert-Schmidt operators $HS(\mathcal{H},\mathcal{E})$,
$\mathrm{vec}(A)\ =\ \sum_{\delta}e_{\delta}\otimes Ae_{\delta}$
for an orthonormal basis $(e_{\delta})_{\delta}$ of $\mathcal{H}$. Define
@@
Let $p=\sum_{w\in\left\langle x\right\rangle_{d+1}}P_{w}\,w$ and $q=\sum_{v\in\left\langle x\right\rangle_{d}}Q_{v}\,v$. Using (5.5) and the standard vectorization identity (see Lemma A.2)
@@
Since $\mathcal{A}_{2d+2}$ is a locally convex topological vector space, by the Hahn–Banach separation theorem (see, e.g., [Dav25, Corollary 3.3.9]), there exist a linear functional $\varphi:\mathcal{A}_{2d+2}\to\mathbb{C},$ and real numbers $\gamma_{1},\gamma_{2}$ such that
@@ 8. Concluding remark and a problem
The reader will have no difficulty verifying that both Theorem 1.1 and Theorem 1.3 remain true if we replace the algebra of coefficients $\mathcal{B}(\mathcal{H})$ with a von Neumann algebra. The proofs carry over to this setting verbatim. On the other hand, we do not know if the results still hold if the coefficients are from a $C^{*}$-algebra.
Appendix A Auxiliary lemmas
For the sake of completeness, this appendix contains two well-known results used in the main body of the paper.
Lemma A.1.
Let $\mathcal{H}$ and $\mathcal{K}$ be Hilbert spaces. Let $(T_{\alpha})_{\alpha}$ be a net in $\mathcal{B}(\mathcal{H})$ that converges to $T\in\mathcal{B}(\mathcal{H})$ in the WOT. Let $X\in\mathcal{B}(\mathcal{K})$ be fixed. Then $(T_{\alpha}\otimes X)_{\alpha}$ converges to $T\otimes X$ in the WOT on $\mathcal{B}(\mathcal{H}\otimes\mathcal{K}).$
Proof.
For $h,h^{\prime}\in\mathcal{H}$ and $k,k^{\prime}\in\mathcal{K}$,
$\langle(T_{\alpha}\otimes X)(h\otimes k),\,h^{\prime}\otimes k^{\prime}\rangle\ =\ \langle T_{\alpha}h,h^{\prime}\rangle\,\langle Xk,k^{\prime}\rangle.$
Since $\langle T_{\alpha}h,h^{\prime}\rangle\to\langle Th,h^{\prime}\rangle$ and $\langle Xk,k^{\prime}\rangle$ is constant, it follows that
$\langle(T_{\alpha}\otimes X)(h\otimes k),h^{\prime}\otimes k^{\prime}\rangle\ \to\ \langle(T\otimes X)(h\otimes k),h^{\prime}\otimes k^{\prime}\rangle.$
By the Uniform Boundedness Principle, we get
$M\ :=\ \sup_{\alpha}\|T_{\alpha}\|<\infty.$
Consequently the family $\{T_{\alpha}\otimes X\}_{\alpha}$ is uniformly bounded in operator norm:
$\|T_{\alpha}\otimes X\|\ \leq\ \|T_{\alpha}\|\ \|X\|\ \leq\ M\|X\|\qquad\text{for all }\alpha.$
Let $\xi,\eta\in\mathcal{H}\otimes\mathcal{K}$ be arbitrary and fix $\varepsilon>0.$ Choose finite sums
$\xi_{0}=\sum_{i=1}^{n}h_{i}\otimes k_{i},\qquad\eta_{0}=\sum_{j=1}^{m}h^{\prime}_{j}\otimes k^{\prime}_{j}$
such that $\|\xi-\xi_{0}\|<\delta$ and $\|\eta-\eta_{0}\|<\delta$, where $\delta>0$ will be chosen below. Write
$\langle(T_{\alpha}\otimes X)\xi,\eta\rangle-\langle(T\otimes X)\xi,\eta\rangle\ =\ A_{\alpha}+B_{\alpha}+C_{\alpha},$
where
$A_{\alpha}\ =$ $\ \langle(T_{\alpha}\otimes X)\xi_{0},\eta_{0}\rangle-\langle(T\otimes X)\xi_{0},\eta_{0}\rangle,$
$B_{\alpha}\ =$ $\ \langle(T_{\alpha}\otimes X)(\xi-\xi_{0}),\eta_{0}\rangle-\langle(T\otimes X)(\xi-\xi_{0}),\eta_{0}\rangle,$
$C_{\alpha}\ =$ $\ \langle(T_{\alpha}\otimes X)\xi,\eta-\eta_{0}\rangle-\langle(T\otimes X)\xi,\eta-\eta_{0}\rangle.$
Note that $A_{\alpha}\to 0$ as $\alpha\to\infty$ because it is a finite linear combination of terms of the form $\langle T_{\alpha}h_{i},h^{\prime}_{j}\rangle\langle Xk_{i},k^{\prime}_{j}\rangle.$ Now
$|B_{\alpha}|\ \leq$ $\ \|T_{\alpha}\otimes X\|\ \|\xi-\xi_{0}\|\ \|\eta_{0}\|+\|T\otimes X\|\ \|\xi-\xi_{0}\|\ \|\eta_{0}\|$
$\ \leq$ $\ 2M\|X\|\,\delta\,\|\eta_{0}\|,$
and
$|C_{\alpha}|\ \leq$ $\ \|T_{\alpha}\otimes X\|\ \|\xi\|\ \|\eta-\eta_{0}\|+\|T\otimes X\|\ \|\xi\|\ \|\eta-\eta_{0}\|$
$\ \leq$ $\ 2M\|X\|\ \|\xi\|\ \delta.$
We can choose $\delta>0$ small enough so that
$|B_{\alpha}|<\varepsilon/3\quad\text{and}\quad|C_{\alpha}|<\varepsilon/3\quad\text{for all $\alpha$}.$
Thus $\langle(T_{\alpha}\otimes X)\xi,\eta\rangle\to\langle(T\otimes X)\xi,\eta\rangle$ for all $\xi,\eta\in\mathcal{H}\otimes\mathcal{K}$, i.e. $T_{\alpha}\otimes X\to T\otimes X$ in the WOT on $B(H\otimes K)$. ∎
The second lemma presents a few basic properties of the vectorization map used in the proof of Proposition 5.3, Subsection 5.5.
Lemma A.2.
Let $A,B\in{HS}(\mathcal{H},\mathcal{E})$ and $P,Q\in B(\mathcal{H})$, $T\in B(\mathcal{E})$.
  1. (1)
    We have
    In particular, $\|\mathrm{vec}(A)\|=\|A\|_{\mathrm{HS}}$, and the value is independent of the choice of the orthonormal basis.
  2. (2)
    The following compatibility with left/right actions holds:
    hence, more generally,
    $(P\otimes T)\,\mathrm{vec}(A)=\mathrm{vec}(T\,A\,P^{*}).$
  3. (3)
    (Inner–product identity used in Proposition 5.3)
Proof.
Fix an orthonormal basis $(e_{r})_{r}$ of $\mathcal{H}$.
(1) Using the definition,
The right-hand side is independent of the choice of orthonormal basis, hence so is the left-hand side.
(2) For the first identity,
(3) Using (2) twice and then (1),

### v2

@@ 5.1. The positive block matrix $S$
For $v,w\in\left\langle x\right\rangle_{d+1}$ define the finite rank operator
@@
Denote by $S$ the $N(d+1)\times N(d+1)$ block operator matrix whose $(v,w)$ entry is $S_{w^{*}v}.$ For $p=\sum_{w\in\langle x\rangle_{d+1}}P_{w}\,w$ and $q=\sum_{v\in\langle x\rangle_{d+1}}Q_{v}\,v$, a direct computation gives the block–trace identity
@@
Define the row operator $P:\bigoplus_{\langle x\rangle_{d+1}}\mathcal{H}\to\mathcal{H}$ by
@@
The positivity of $S$ can be concluded exactly as in the proof of Lemma 4.4. Because $J$ and $\langle x\rangle_{d+1}$ are finite, $S$ is finite rank.
@@ 5.2. An Auxiliary Hilbert space $\mathcal{M}$
We construct a finite-dimensional Hilbert space $\mathcal{M}$ from the psd sesquilinear form induced by the psd operator $S.$ The Hilbert space $\mathcal{E}$ is constructed as a subspace of $\mathcal{M}$ in Subsection 5.4.
@@
$\mathcal{N}\ =\ \{z\in V:\langle z,z\rangle_{V}=0\}$
@@
$\mathcal{M}\ =\ V/\mathcal{N}.$
Since $S$ has finite rank, $\mathcal{M}$ is finite dimensional.
@@ 5.3. The coordinate maps $\Phi(w)$
For each $w\in\langle x\rangle_{d+1}$ define $\Phi(w):\mathcal{H}\to\mathcal{M}$ by
$\Phi(w)\xi\ =\ \big[(\delta_{u,w}\xi)_{u\in\langle x\rangle_{d+1}}\,\big],$
@@ 5.4. The Hilbert space $\mathcal{E}$ and the operator tuple $Y$
$L_{x_{i}}\big(\Phi(w)\xi\big)\ =\ \Phi(x_{i}w)\,\xi,\qquad(w\in\left\langle x\right\rangle_{d},\ \xi\in\mathcal{H}).$ (5.4)
@@ 5.4.1. The $L_{x_{j}}$ are well-defined
Since the vectors $\Phi(v)\eta$ span $\mathcal{M}$, we have $L_{x_{i}}z\in\mathcal{N},$ proving that $L_{x_{i}}$ is well-defined.
Let $P_{\mathcal{E}}$ be the orthogonal projection of $\mathcal{M}$ onto $\mathcal{E}.$ For $i=1,\dots,g,$ define $Y_{i}:\mathcal{E}\to\mathcal{E}$ by
@@ 5.5. The representing vector and evaluation
We construct the representing vector $\gamma$ and complete the proof. It is here that the vectorization map of Lemma 5.3 appears.
Define
@@
Let $p=\sum_{w\in\left\langle x\right\rangle_{d+1}}P_{w}\,w$ and $q=\sum_{v\in\left\langle x\right\rangle_{d}}Q_{v}\,v$. Using (5.5) and the standard vectorization identity (see Lemma 5.3)
@@
Since $\mathcal{A}_{2d+2}$ is a locally convex topological vector space and $\mathcal{C}_{d+1}$ is closed by Proposition 4.1, the Hahn–Banach separation theorem (see, e.g., [Dav25, Corollary 3.3.9]) implies that there exist a WOT continuous linear functional $\varphi:\mathcal{A}_{2d+2}\to\mathbb{C},$ and real numbers $\gamma_{1},\gamma_{2}$ such that
@@ 8. Concluding remark and a problem
The reader will have no difficulty verifying that both Theorem 1.1 and Theorem 1.3 remain true if we replace the algebra of coefficients $\mathcal{B}(\mathcal{H})$ with a von Neumann algebra. The proofs carry over to this setting in a straightforward way. On the other hand, we do not know if the results still hold if the coefficients are from a $C^{*}$-algebra.